For 46,400 Euros you can have a new Ferrari with a V-12 engine rated at 808...

  1. For 46,400 Euros you can have a new Ferrari with a V-12 engine rated at 808 hp. Your new car has a maximum speed of 340 km/hr at 6,300 rpm.

A. Determine its maximum rimpull when operated at maximum speed if its engine efficiency is 90%. The Ferrari weighs 4,155 lb and is operated over a concrete road with rolling resistance of 40 lb/ton and a slope of 3%.

B. Determine the maximum external pull the Ferrari has when operated in sixth gear at maximum speed.
Answer #1

Given

Engine power P = 808 hp , Maximum speed = V= 340 Km/hr = 211.2662 mph, Engine efficiency = 90%, Ferrari weight = 4155lb , Rolling resistance = 40lb/ton, slope = 3%

Maximum rimpull is given as

Maximum external pull the ferrari has = Maximum Rim pull - ( Pull required to overcome grade + Pull required to overcome rolling resistance)

{Pull required to overcome grade = Weight X slope = 4155 X 0.03 = 124.65 lb

Pull required to overcome rolling resistance = weight in ton X rolling resistance = (4155/2000) tons X 40 = 83.1 lb}

Maximum external pull = 1290.788 -(124.65+83.1)

Maximum external pull = 1083.038 lb

Note - If we want to consider the drag resistance due to air than it is also deducted from maximum rim pull for determining the external pull